OpenStudy (anonymous):

Graph the equation. 3x+5y=11

OpenStudy (anonymous):

First you need to get it into 'y=mx+b' form so solve for y

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

Yeah, but what did you get?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

3x+5y=11 subtract 3x 5y=-3x+11 Now divide by 5 what did you get?

OpenStudy (anonymous):

y=3/5 + 11/5

OpenStudy (anonymous):

Right, but you left out the x so, y = 3/5x + 11/5

OpenStudy (anonymous):

Now that it is in y=mx + b form, m being the slope and b being the y intercept you can graph it

OpenStudy (anonymous):

do you know how to graph this?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Well 11/5 is a little more than 2, so the y intercept is going to be a little more than 2. The slope is -3/5 Slope - rise/run - Since it's -3 it's going to go down 3, and right 5

OpenStudy (anonymous):

 It's going to look roughly something like that